OpenStudy (anonymous):

HOW WOULD I DO THE FOLLOWING QUESTION: find all the values of c for which the following exists...

OpenStudy (anonymous):

\[\lim_{x \rightarrow c} \frac{ x^3+3x^2-4x-12 }{x-c}\]

OpenStudy (anonymous):

the limit will only exist if you can cancel the \(x-c\) top and bottom, otherwise you will have a zero in the denominator, and not a zero in the numerator this is another way of saying that \(c\) must be a zero of \[x^3+3x^2-4x-12\] because if so, then by the factor theorem you can factor \[x^3+3x^2-1x-12\] as \((x-c)(something)\)

OpenStudy (anonymous):

in other words, you solve this question by finding the zeros of the numerator those are your possible values for \(c\)

OpenStudy (anonymous):

am i correct when i say that c cannot equal 2 or -2 or -3 as the limit does not exist

OpenStudy (anonymous):

you are asked for values of c that work, not values of c that do not work

OpenStudy (anonymous):

if those are the zeros of the numerator, then those are the only values of c that DO work i.e. the only values of c for which the limit exists otherwise it will not exist

OpenStudy (anonymous):

but none of those values (they are in the numerator) cancel with the denominator.. doesn't the denominator need to cancel in order for the limit to exist?

OpenStudy (anonymous):

yes, the question asks which ones give a limit, not which ones do not

OpenStudy (anonymous):

set the numerator equal to zero and solve those are the numbers that will give a limit any other numbers will not

OpenStudy (anonymous):

so all i have to write (this is an assignment)would be.. \[\lim_{x \rightarrow c}\frac{ x^3+3x^2-4x-12 }{x-c}\] \[x^3+3x^2-4x-12 = 0\] \[x^2(x+3)-4(x+3)\] \[(x^2-4)(x+3)\] \[(x+2)(x-2)(x+3)\] x=-2, x=2, or x=3

OpenStudy (anonymous):

ok so those are the numbers for which you WOULD get a limit

OpenStudy (anonymous):

how though? if i plug it in i get 0/0

OpenStudy (anonymous):

ok lets go slow also note that if \(x+3\) is a factor, then the zero is \(-3\)

OpenStudy (anonymous):

\[\frac{0}{0}\] is your friend it means you could get a limit \[\frac{5}{0}\] means no way you can get a limit

OpenStudy (anonymous):

nevermind i got real numbers

OpenStudy (anonymous):

so if the numerators is not zero, but the denominator is, then no limit possible

OpenStudy (anonymous):

thank you but the way i had it done out here, will that be ok to write out on my assignment or should it written out differently?

OpenStudy (anonymous):

you should say that those number WILL give you a limit, and only those numbers lets compute one \[\lim_{x\to 2}\frac{(x+2)(x-2)(x+3)}{x-2}=\lim_{x\to 2}(x+2)(x+3)=4\times 5=20\]

OpenStudy (anonymous):

Thanks again! don't know what i would've done without your help tonight!!!

OpenStudy (anonymous):

yw, hope this one is more clear now